# Interacting tachyon with varying mass dark matter

### v1

@@ V Comparison of the model with the observational data
To check the reliability of our model further, the predicted evolution of the Hubble parameter and distance modulus have also been studied and compared that with the observational Hubble parameter and Type Ia Supernova (SNIa) datasets in figure 8. In figure 8, we have plotted the nature of the Hubble parameter as a function of redshift $z$ and compared it with that of the Hubble data with $1\sigma$ error bars obtained from the compilation of 31 points of $H(z)$ measurements [58]. Also, the theoretical evolution of distance modulus, which is the difference between the apparent and absolute magnitudes of the observed SNIa, as a function of $z$ using this model is compared with the SNIa data points [59] and is shown in figure 8. It is clear from figures 8 and 8 that the theoretical model show proximity with observational results at low redshift very well.
@@ VI Concluding Remarks
In this paper, we have studied a cosmological model of interacting tachyon with varying-mass dark matter particles in the background of FLRW universe where the tachyonic fluid plays the role of DE and pressureless dust as dark matter. In this variable mass particle (VAMP) mechanism, it is assumed that the mass of the DM particles is time dependent through the scalar field $\phi$ in the sense that the decaying of DM particles reproduces the scalar field. We have studied the cosmological model in the context of dynamical analysis by transforming suitable dimensionless variables. We considered the mass of DM particles and the scalar field potential to be varied with exponential function of $\phi$. In this study we considered the exponential mass dependence and exponential potential function of scalar field $\phi$, i.e., $Q_{m}(\phi)=Q_{m0}~{}\mbox{exp}\{-\alpha H_{0}\phi\}$ and $V(\phi)=V_{0}~{}exp\{\beta H_{0}\phi\}$, where $Q_{m0}$ and $V_{0}$ are constant and $\alpha$ and $\beta$ are constant parameters. As a result, the autonomous system is reduced to be a 3D system. We have studied the phase space analysis of the system (24). We have extracted seven critical points from this system. The critical points and the cosmological parameters are presented in the table 1 and the eigenvalues of the Jacobian matrix are displayed in the table 2. We observed that all the critical points are non-hyperbolic in nature. Among them, the set of critical points $A$, $D$ and $E$ are non-isolated sets having all the eigenvalues zero for which it is very hard to find the nature of these sets. However, numerical investigations are carried out to find the nature of these sets. These sets behave as dust dominated decelerated intermediate phase of the universe (since $\Omega_{m}=1$$\omega_{eff}=0,~{}q=\frac{1}{2}$). These solutions are physically interested because they can successfully describe the phase of recent past evolution of the universe. On the other hand, the DE dominated solutions namely, the points $B$ and $C$ are non-hyperbolic having the values of cosmological parameters $\Omega_{\phi}=1,~{}\omega_{eff}=-1,~{}q=-1$ can show the de Sitter expansion of the universe. Due to physical importance, we have performed the center manifold theorem to study their stability. We note that the character of the flow on the center manifold does not depend on the parameter $\beta$ (see Equations (57) and (68)). For the case of critical points $B$ and $C$, the flow on the center manifold is topologically equivalent to $\frac{dZ}{dN}\approx\frac{3}{2}Z^{3}$. It is to be noted that the vector field on the center manifold is repelling away from the critical points. On the other hand, the vector field along the eigenvector of the eigenvalue $\lambda_{2}=-3$ is attracting towards the critical points. So the late-time solutions near the critical points $B$ and $C$ depend on initial conditions [60]. Finally, we obtained the DE-DM scaling solutions described by the critical points $F$ and $G$ which are also non-hyperbolic points in the phase space and due to complicated nature we have performed numerical investigation in fig. (4) to show their stability. We can conclude that the points can show the late time scaling attractors solving the coincidence problem. If the trajectory of the vector field in the phase-space starts in the vicinity of the critical point $F$ or $G$, it goes towards $z=1$ plane in a spiral-like manner (see fig (4)), and the values of cosmological parameters oscillate. The nearer the trajectory goes towards of $z=1$ plane the amplitude of oscillation decreases and the values of deceleration parameter oscillate between $-0.18$ to $-0.38$ (approx) whereas the values of effective equation of state parameter oscillate between $-0.4$ to $-0.6$ (approx). On the other hand, the density parameters oscillate between $0.4$ to $0.6$ (a scaling solution) near the $z=1$ plane (see fig 5). But, immediately after, the trajectory touches $z=1$ plane the amplitudes of oscillation of the cosmological parameters increase rapidly as the critical points $F$ and $G$ are unstable focus in nature. Thus, the strange cosmological behavior is to be observed in the vicinity of $z=1$ plane. The nature of solutions bifurcates from scalar field dominated to matter dominated by passing through the scaling solution and deceleration to acceleration era of expansion by passing through the quintessence boundary and the other way round. So, in our model, $z=1$ plane turns out to be a strange attractor in the perspective of strange cosmological behavior corresponding to the vector field near the plane.
@@
We have also investigated the evolution of cosmographical parameters with the help of statefinder pair diagnostic tool and examined the reliability of our model by comparing with the observational data such as Hubble parameter and SNIa. From the statefinder diagnostic analysis, it has been found that the model explains the late-time cosmic acceleration and also transits from early decelerating phase to an accelerating phase. It is also observed that the model shows proximity with observational data very well.
@@ Acknowledgments
The author Goutam Mandal acknowledges UGC, Government of India for providing Senior Research Fellowship [Award Letter No. F.82-1/2018(SA-III)] for Ph.D. The authors would like to thank Md. Arif Shaikh for helpful discussions.

### v2

@@ V Comparison of the model with the observational data
To check the reliability of our model further, the predicted evolution of the Hubble parameter and distance modulus have also been studied and compared that with the observational Hubble parameter and Type Ia Supernova (SNIa) datasets in figure 9. In figure 9, we have plotted the nature of the Hubble parameter as a function of redshift $z$ and compared it with that of the Hubble data with $1\sigma$ error bars obtained from the compilation of 31 points of $H(z)$ measurements [79]. Also, the theoretical evolution of distance modulus, which is the difference between the apparent and absolute magnitudes of the observed SNIa, as a function of $z$ using this model is compared with the SNIa data points [80] and is shown in figure 9. It is clear from figures 9 and 9 that the theoretical model show proximity with observational results at low redshift very well.
@@ VI Concluding Remarks
In this paper, we have studied a cosmological model of interacting tachyon with varying-mass dark matter particles in the background of FLRW universe where the tachyonic fluid plays the role of DE and pressureless dust as dark matter. In this variable mass particle (VAMP) mechanism, it is assumed that the mass of the DM particles is time dependent through the scalar field $\phi$ in the sense that the decaying of DM particles reproduces the scalar field. We have studied the cosmological model in the context of dynamical analysis by transforming suitable dimensionless variables. We considered the mass of DM particles and the scalar field potential to be varied with exponential function of $\phi$. In this study we considered the exponential mass dependence and exponential potential function of scalar field $\phi$, i.e., $Q_{m}(\phi)=Q_{m0}~{}\mbox{exp}\{-\alpha H_{0}\phi\}$ and $V(\phi)=V_{0}~{}exp\{\beta H_{0}\phi\}$, where $Q_{m0}$ and $V_{0}$ are constant and $\alpha$ and $\beta$ are constant parameters. As a result, the autonomous system is reduced to be a 3D system. We have studied the phase space analysis of the system (LABEL:autonomous_system_2). We have extracted seven critical points from this system. The critical points and the cosmological parameters are presented in the table 1 and the eigenvalues of the Jacobian matrix are displayed in the table 2. We observed that all the critical points are non-hyperbolic in nature. Among them, the set of critical points $A$, $D$ and $E$ are non-isolated sets having all the eigenvalues zero for which it is very hard to find the nature of these sets. However, numerical investigations are carried out to find the nature of these sets. These sets behave as dust dominated decelerated intermediate phase of the universe (since $\Omega_{m}=1$$\omega_{eff}=0,~{}q=\frac{1}{2}$). These solutions are physically interested because they can successfully describe the phase of recent past evolution of the universe. On the other hand, the DE dominated solutions namely, the points $B$ and $C$ are non-hyperbolic having the values of cosmological parameters $\Omega_{\phi}=1,~{}\omega_{eff}=-1,~{}q=-1$ can show the de Sitter expansion of the universe. Due to physical importance, we have performed the center manifold theorem to study their stability. We note that the character of the flow on the center manifold does not depend on the parameter $\beta$ (see eqs. (44) and (54)). For the case of critical points $B$ and $C$, the flow on the center manifold is topologically equivalent to $\frac{dZ}{dN}\approx\frac{3}{2}Z^{3}$. It is to be noted that the vector field on the center manifold is repelling away from the critical points. On the other hand, the vector field along the eigenvector of the eigenvalue $\lambda_{2}=-3$ is attracting towards the critical points. So the late-time solutions near the critical points $B$ and $C$ depend on initial conditions [81]. Finally, we obtained the DE-DM scaling solutions described by the critical points $F$ and $G$ which are also non-hyperbolic points in the phase space and due to complicated nature we have performed numerical investigation in figure 5 to show their stability. We can conclude that the points can show the late time scaling attractors solving the coincidence problem. If the trajectory of the vector field in the phase-space starts in the vicinity of the critical point $F$ or $G$, it goes towards $z=1$ plane in a spiral-like manner (see figure 5), and the values of cosmological parameters oscillate. The nearer the trajectory goes towards of $z=1$ plane the amplitude of oscillation decreases and the values of deceleration parameter oscillate between $-0.18$ to $-0.38$ (approx) whereas the values of effective equation of state parameter oscillate between $-0.4$ to $-0.6$ (approx). On the other hand, the density parameters oscillate between $0.4$ to $0.6$ (a scaling solution) near the $z=1$ plane (see figure 6). But, immediately after, the trajectory touches $z=1$ plane the amplitudes of oscillation of the cosmological parameters increase rapidly as the critical points $F$ and $G$ are unstable focus in nature. Thus, the strange cosmological behavior is to be observed in the vicinity of $z=1$ plane. The nature of solutions bifurcates from scalar field dominated to matter dominated by passing through the scaling solution and deceleration to acceleration era of expansion by passing through the quintessence boundary and the other way round. So, in our model, $z=1$ plane turns out to be a strange attractor in the perspective of strange cosmological behavior corresponding to the vector field near the plane.
@@
We have also investigated the evolution of cosmographical parameters with the help of statefinder pair diagnostic tool and examined the reliability of our model by comparing with the observational data such as Hubble parameter and SNIa. From the statefinder diagnostic analysis, it has been found that the model explains the late-time cosmic acceleration and also transits from early decelerating phase to an accelerating phase. It is also observed that the model shows proximity with observational data very well. The tachyon model presents a compelling alternative to conventional cosmological scenarios, offering richer and more complex dynamics. Its potential to describe certain aspects of the universe warrants further attention. A thorough statistical evaluation, using data from Type Ia Supernovae, Baryon Acoustic Oscillations, the Cosmic Microwave Background, and measurements of large-scale structure (such as $\sigma_{8}$), is essential to test its viability and refine its parameters. While this paper lays the groundwork, such comprehensive analysis is beyond its current sope and is reserved for future research efforts.
@@ Acknowledgments
The authors would like to thank the anonymous referees for their critical review and suggestions as a result of which the paper has improved significantly. The author Goutam Mandal acknowledges UGC, Government of India for providing Senior Research Fellowship [Award Letter No. F.82-1/2018(SA-III)] for Ph.D. The authors would like to thank Md. Arif Shaikh for helpful discussions. The authors are also grateful to Dr. Supriya Pan for helpful discussion on interaction in dark sectors. SKB would like to acknowledge Inter-University Centre for Astronomy and Astrophysics (IUCAA), India for providing research facilities.